OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − 8 = x2 − 6x?

OpenStudy (anonymous):

x intercept means, it crosses X-axis i.e., y=0 substitute y=0 in the equation and solve for "x" there will be 2 values.

OpenStudy (johnweldon1993):

exactly...so after doing as what @electrokid has said 0 - 8 = x² - 6x -8 = x² - 6x add 8 to both sides x² - 6x + 8 = 0 now factor this...what 2 numbers multiply to equal 8...and at the same time add to become -6? 4 and 2 sound like a good fit (x - 4)(x - 2) lets check it x² -2x -4x +8 x² - 6x + 8 indeed ...so ...using what we now have (x - 4)(x - 2) what makes these = 0? ? - 4 = 0? ? - 2 = 0?